# Markov Perfect Equilibria in

### v1

@@ 4.2 The mean-field Nash-Lasry-Lions equation
We now discuss the (MF-NLL) equation and first recall that any solution to the (MF-NLL) equation, regardless of its regularity in the $\mu$ variable, defines a solution to the MFG system for any initial condition $(t,\mu)\in{\mathcal{T}}\times\Sigma^{d-1}$, and hence gives rise to solutions to the mean-field game, see Proposition 2.18 (ii).
@@
Let Assumptions 2.5 and 2.7 hold. Then, for any Lipschitz continuous function $\varphi:\mathscr{X}\to\mathbb{R}$, there exists a unique map $\hat{\alpha}(x,\mu):\mathscr{X}\to\Sigma^{d-1}$ that satisfies
@@
where $\gamma$ and $L_{\partial\ell}$ are constants from Assumptions 2.7 and 2.19, respectively.
@@
${\mathcal{S}}:=\{\alpha:{\mathcal{X}}\to\Sigma^{d-1}\,:\,\alpha_{y}(x)=0\ \forall y\notin S(x)\},\quad|\alpha|_{\mathcal{S}}:=\max_{x\in{\mathcal{X}}}\sum_{y\in S(x)}|\alpha_{y}(x)|,\quad\alpha\in{\mathcal{S}}.$
@@
We first note that, if $\mu\in\Sigma^{d-1}$ is a row vector and $A\in\mathbb{R}^{d\times d}$, then $|\mu\cdot A|_{1}\leq\max_{1\leq i\leq d}|A_{i,:}|_{1}$. Together with Lemma 3.3,
$|\Phi_{x}(\alpha)-\Phi_{x}(\tilde{\alpha})|_{1}\leq\frac{h}{\gamma}\sum_{y\in S(x)}|\varphi(y,\mu\cdot\alpha)-\varphi(y,\mu\cdot\tilde{\alpha})|\leq\frac{hL_{\varphi}}{\gamma}\,|\mu\cdot(\alpha-\tilde{\alpha})|_{1}\leq\frac{hL_{\varphi}}{\gamma}\,\max_{x\in{\mathcal{X}}}|\alpha(x)-\tilde{\alpha}(x)|_{1},$
so that $\Phi(\cdot)$ is a contraction if $h<\gamma/L_{\varphi}$ where $L_{\varphi}=m\max_{x\in{\mathcal{X}}}\mathrm{Lip}_{1}(\varphi(x,\cdot))$. By Banach’s fixed-point theorem, there exists a unique fixed point in ${\mathcal{S}}$. Since $\mu\in\Sigma^{d-1}$ was arbitrary, this construction yields a unique fixed point $\hat{\alpha}(x,\mu)$. The Lipschitz constant of $\hat{\alpha}$ can be deduced similarly to the finite-player case, see the proof of Proposition 3.4. ∎
@@ 4.3 One-step example
${\mathcal{X}}=\{0,1\},\quad\tilde{A}(h)=[\sigma^{2}h,1-\sigma^{2}h],\quad\ell(x,a)=\frac{1}{2h}a^{2},\quad g(x,\mu)=\mu\,\chi_{x=0}+(1-\mu)\,\chi_{x=1},$
for $(x,a,\mu)\in{\mathcal{X}}\times\tilde{A}(h)\times[0,1]$. As in the two-player Example 3.6, addition on ${\mathcal{X}}$ is defined modulo $2$ and we identify controls $\alpha(x)=(\alpha_{0}(x),\alpha_{1}(x))$ with the probability $\alpha_{x+1}(x)\in\tilde{A}(h)$ of changing the current state. Similarly, we only keep track of the fraction $\mu\in[0,1]$ of the population in state $1$. Given a flow $\mu=(\mu_{0},\mu_{h})\in[0,1]^{2}$ of probabilities representing the fractions of players in state $1$, a typical player solves
@@
for every $x\in{\mathcal{X}}$. Here, $\mathbb{P}^{\alpha}[X_{h}=x+1\,|\,X_{0}=x]=\alpha(x)$, $x\in{\mathcal{X}}$. The cost functional is chosen in a way that leads to synchronization with the population. Given the initial distribution $\mu_{0}$, mean-field game equilibria are characterized by the following system:
$\begin{cases}\alpha(x)=\underset{a\in A(h)}{\mathrm{arg\,min}}\,\{a^{2}/(2h)+(1-a)g(x,\mu_{h})+ag(x+1,\mu_{h})\},\\ \mu_{h}=\alpha(0)(1-\mu_{0})+(1-\alpha(1))\mu_{0}.\end{cases}$
@@
with $\alpha^{\mu_{h}}(0):=[h(2\mu_{h}-1)]\lor(\sigma^{2}h)\land(1-\sigma^{2}h)$ and $\alpha^{\mu_{h}}(1):=[h(1-2\mu_{h})]\lor(\sigma^{2}h)\land(1-\sigma^{2}h)$. This is a fixed-point equation for the Nash equilibrium distribution $\mu_{h}\in[0,1]$. If $\sigma^{2}=0$, then there are three equilibria starting from the uniform distribution become if $h\geq 1$. If $\sigma^{2}>0$, we have the following result, which follows by inspection of the above fixed-point equation, see Figure 1 for an illustration.
@@
    If $h\in\{h_{*},h^{*}\}$, then there are exactly three NE.
@@
    If $h\in(h_{*},h^{*})$, then there are exactly five NE.
@@
Figure 1: Nash equilibria starting from $\mu_{0}=1/2$ when $\sigma^{2}>0$.
@@ 5 Short-time results
Theorem 2.10 establishes the existence and uniqueness of the solution to ($N$-NLL) for a time step $h<h_{*}(N,T)$ that depends on both the number of players and the time horizon of the game. In fact, as stated in Theorem 2.20, there exist time steps that guarantee uniqueness for any finite-player game as well as for the mean-field limit. We proceed to establish this result.
@@
    (Finite-player case). Let $\varphi:\mathscr{X}_{N}\to\mathbb{R}$ be given. Let $\hat{\alpha}$ be as in Proposition 3.4, and define
    $\psi(x,z)=H(x,z,E(x,z,\varphi,\hat{\alpha}),\hat{\alpha}(x,z)),\quad(x,z)\in\mathscr{X}_{N}.$
@@
    (Mean-field case). Let $\varphi:\mathscr{X}\to\mathbb{R}$ be given. Let $\hat{\alpha}$ be as in Proposition 4.1, and define
    $\psi(x,\mu)=H(x,\mu,\varphi(\cdot,\mu\cdot\hat{\alpha}(\cdot,\mu)),\hat{\alpha}(x,z)),\quad(x,\mu)\in\mathscr{X}.$
@@
$|\psi(x,z)-\psi(x,\tilde{z})|$ $\leq\sup_{a\in A(h,x)}|(\ell(x,z,a/h)-\ell(x,\tilde{z},a/h))h+a\cdot(E(x,z,\varphi,\hat{\alpha})-E(x,\tilde{z},\varphi,\hat{\alpha}))|$
@@
(b) For $\mu,\tilde{\mu}\in\Sigma^{d-1}$,
$|a\cdot(\varphi(\cdot,\mu\cdot\hat{\alpha}(\cdot,\mu))-\varphi(\cdot,\tilde{\mu}\cdot\hat{\alpha}(\cdot,\tilde{\mu})))|$ $\leq L_{\varphi}/m\,|\mu\cdot\hat{\alpha}(\cdot,\mu)-\tilde{\mu}\cdot\hat{\alpha}(\cdot,\tilde{\mu})|_{1}$
@@
$|\psi(x,\mu)-\psi(x,\tilde{\mu})|$ $\leq\sup_{a\in A(h,x)}|(\ell(x,\mu,a/h)-\ell(x,\tilde{\mu},a/h))h+a\cdot(\varphi(\cdot,\mu\cdot\hat{\alpha}(\cdot,\mu))-\varphi(\cdot,\tilde{\mu}\cdot\hat{\alpha}(\cdot,\tilde{\mu})))|$
@@
Let Assumptions 2.5, 2.7 and 2.19 hold. There exist constants $h_{*},T_{*}>0$ depending only on the cost functions $\ell,g$ and $m=m(d)$, such that, for all $h>0$ and $K\geq 1$, the following holds: If $h<h_{*}$ and $Kh<T_{*}$, then there exists a unique solution $v^{\infty}$ to (MF-NLL) and a unique solution $(v^{(N)},\alpha^{(N)})$ to ($N$-NLL) on $\bar{\mathcal{T}}=\{0,\ldots,Kh\}$ for every $N\geq 1$. Furthermore,
$\sup_{N\in\mathbb{N}\cup\{\infty\}}\max_{(t,x)\in\bar{\mathcal{T}}\times{\mathcal{X}}}\,[\mathrm{Lip}_{1}(v^{(N)}(t,x,\cdot))+\mathrm{Lip}_{1}(\alpha^{(N)}(t,x,\cdot))]<\infty.$
@@
$L(K)$ $=L(0)+\sum_{k=1}^{K}(L(k)-L(k-1))\leq L(0)+h\sum_{k=1}^{K}\Bigl{(}mL_{\ell}+L(k-1)\frac{L_{\partial\ell}+L(k-1)}{\gamma-hL(k-1)}\Bigr{)}\leq L(0)+Kh\,\tilde{M}\leq M,$
@@
Step 2. In light of the previous step, we may choose any $M>L(0)$ and define $h_{*}:=\gamma/M$ and $T_{*}:=(M-L(0))/\tilde{M}$. Note that, in particular, we have $h<\gamma/L(k)$ for all $k=0,\ldots,K$. By Propositions 3.4 and 4.1, we can recursively solve the equations ($N$-NLL) and (MF-NLL) on $\{0,\ldots,Kh\}$. Finally, note that the Lipschitz constants are bounded by $M$. ∎
Remark 5.3.
If the number of admissible transitions $m=m(d)$ is independent of $d$, then so are $h_{*}$ and $T_{*}$. This is, for instance, the case in Example 2.3.
@@ 6 Mean-field convergence
$\mathbb{E}[\varphi^{(N)}(x,Z^{\prime}(x,z^{(N)},\alpha^{(N)}))]\to\varphi^{(\infty)}(x,\sum_{y\in{\mathcal{X}}}\mu(y)\alpha^{(\infty)}(y,\mu)),\quad\text{as}\quad N\to\infty.$
@@
Step 1. Let $h_{*},T_{*}$ be as in Proposition 5.2. Clearly, the solutions $(v^{(N)},\alpha^{(N)})_{N\geq 1}$ of ($N$-NLL) form a uniformly bounded equicontinuous sequence of continuous functions from $\Sigma^{d-1}$ to $[0,\infty)^{\bar{\mathcal{T}}\times{\mathcal{X}}}\times(\Sigma^{d-1})^{{\mathcal{T}}\times{\mathcal{X}}}$. Here, we linearly interpolated in the $z$-variable. Hence, by the Arzelà-Ascoli theorem, there exists a $(\tilde{v}^{(\infty)},\tilde{\alpha}^{(\infty)})$ such that, along a subsequence, $(v^{(N)},\alpha^{(N)})\to(\tilde{v}^{(\infty)},\tilde{\alpha}^{(\infty)})$ uniformly as $N\to\infty$. We replace the original sequence by this subsequence.
@@
$\tilde{\alpha}^{(\infty)}(t,x,\mu)=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\,H(x,\mu,\tilde{\varphi}^{(\infty)}(\mu^{T}\tilde{\alpha}^{(\infty)}(\cdot,\mu)),a),\quad(t,x,\mu)\in{\mathcal{T}}\times\mathscr{X}.$
@@
For a given $\mu\in\Sigma^{d-1}$, choose $z^{(N)}\in\mathscr{X}_{N}$ such that $z^{(N)}\to\mu$ as $N\to\infty$. By Lemma 6.1,
$H(x,\mu,\tilde{v}^{(\infty)}(t+h,\mu\cdot\tilde{\alpha}^{(\infty)}(t,\cdot,\mu)),\tilde{\alpha}^{(\infty)}(t,x,\mu))\leq H(x,\mu,\tilde{v}^{(\infty)}(t+h,\mu\cdot\tilde{\alpha}^{(\infty)}(t,\cdot,\mu)),a)$
for $(t,x,\mu,a)\in{\mathcal{T}}\times\mathscr{X}\times A(h,x)$. Hence, $(\tilde{v}^{(\infty)},\tilde{\alpha}^{(\infty)})$ is a solution of (MF-NLL) on $\bar{\mathcal{T}}$, which has a unique solution $(V,\alpha)$. This argument shows that any subsequence of $(v^{(N)},\alpha^{(N)})_{N\geq 1}$ has a further subsequence that converges to $(V,\alpha)$ uniformly, hence the entire sequence converges to $(V,\alpha)$ uniformly. ∎
@@ 7 Continuous-time limit
This section proves Theorem 2.23. Throughout this section we assume
$A(h)=A(h,x)=\{a\in\Sigma^{d-1}:a_{y}\geq h\sigma^{2}\ \forall y\in{\mathcal{X}}\},\quad\text{for some fixed}\ \ \sigma^{2}\geq 0.$
@@
Let $(\alpha^{(K)},\alpha^{(\infty)}):\mathscr{X}_{N}\to\Sigma^{d-1}\times\mathbb{R}^{d}$, $K\geq 1$, be functions such that $\mathfrak{a}^{(K)}/h^{(K)}\to\alpha^{(\infty)}$ as $K\to\infty$ where $\mathfrak{a}^{(K)}(x,z):=\alpha^{(K)}(x,z)-e_{x}$ for $(x,z)\in[0,T]\times\mathscr{X}_{N}$. Then, for all $(x,z,z^{\prime})\in\mathscr{X}_{N}\times\Sigma^{d-1}_{N}$,
@@
and, more precisely, $\lim_{K\to\infty}\mathbb{P}[Z^{\prime}(x,z,a^{(K)})=z]=1$.
@@
Fix $(x,z)\in\mathscr{X}_{N}$. For any state $y\in{\mathcal{X}}$ and any empirical distribution $s\in\Sigma^{d-1}_{Nz_{y}}$ of untagged players in state $y$,
$\mathbb{P}[Z^{\prime}(x,y,z,\alpha^{(K)})=s]=\frac{(Nz_{y})!}{(Ns_{1})!\cdots(Ns_{d})!}\,[1+\mathfrak{a}_{y}^{(K)}(y,z+e_{xy})]^{Ns_{y}}\prod_{i\neq y}[\mathfrak{a}^{(K)}_{i}(y,z+e_{xy})]^{Ns_{i}}.$
By inspection of this formula we deduce
@@
    if $s=z_{y}e_{y}$, i.e., all untagged players in state $y$ remain in $y$ after $h^{(K)}$ amount of time, then
    $\mathbb{P}[Z^{\prime}(x,y,z,\alpha^{(K)})=z_{y}e_{y}]=[1+\mathfrak{a}^{(K)}_{y}(y,z+e_{xy})]^{Nz_{y}}\to 1,\quad\text{as}\ K\to\infty,$
    using $\mathfrak{a}^{(K)}_{y}(t,y,z+e_{xy})\to 0$;
@@
    if there exists $w\in{\mathcal{X}}\setminus\{y\}$ such that $s=z_{y}e_{y}+e_{wy}$, i.e., exactly one untagged player in state $y$ changes to a state $w$, then
    $\frac{1}{h^{(K)}}\mathbb{P}[Z^{\prime}(x,y,z,\alpha^{(K)})=z_{y}e_{y}+e_{wy}]=\frac{(Nz_{y})!}{(Nz_{y}-1)!}\,[1+\mathfrak{a}_{y}^{(K)}(y,z+e_{xy})]^{Nz_{y}-1}\frac{\mathfrak{a}^{(K)}_{w}(y,z+e_{xy})}{h^{(K)}},$
    which converges to $Nz_{y}\,\alpha^{(\infty)}_{w}(t,y,z+e_{xy})$;
@@
    in all other cases,
    $\frac{1}{h^{(K)}}\mathbb{P}[Z^{\prime}(x,y,z,\alpha^{(K)})=s]=\frac{(Nz_{y})!}{(Ns_{1})!\cdots(Ns_{d})!}\,[1+\mathfrak{a}_{y}^{(K)}(y,z+e_{xy})]^{Ns_{y}}\frac{1}{h^{(K)}}\prod_{i\neq y}[\mathfrak{a}^{(K)}_{i}(y,z+e_{xy})]^{Ns_{i}}$
    converges to zero as $\mathfrak{a}_{i}^{(K)}(t,y,z+e_{xy})/h^{(K)}$ stays bounded as $K\to\infty$ for all $i$ and there are at least two states $i\neq y$ present in the product.
Next, let $z^{\prime}\neq z$ in $\Sigma^{d-1}_{N}$ be given and write $z^{\prime}=s(1)+\cdots+s(d)$ for some $s(y)\in\Sigma^{d-1}_{Nz_{y}}$, with $\Sigma^{d-1}_{0}:=\{0_{\mathbb{R}^{d}}\}$. We let $\Sigma(z^{\prime})$ denote the set of all such $s(1),\ldots,s(d)$. Then, the claim follows from the previous observation, together with the explicit formula
@@
Now, for each $K\geq 1$, set $h^{(K)}:=T/K$ and define the discretization $0=s_{0}^{(K)}<s_{1}^{(K)}<\cdots<s_{k}^{(K)}=T$ of the time interval $[0,T]$ by $s^{(K)}_{k}:=kT/K$. For each $K\geq 1$, let $(v^{(K)},\alpha^{(K)})$ solve the ($N$-NLL) equation, which is uniquely defined for sufficiently large $K$, say $K\geq K_{0}$. We extend these functions to $[0,T]\times\mathscr{X}_{N}$ by linear interpolation in the time variable and set
$\mathfrak{a}^{(K)}(t,x,z):=\alpha^{(K)}(t,x,z)-e_{x},\quad(t,x,z)\in[0,T]\times\mathscr{X}_{N}.$
In the continuous-time problem, let $v_{\mathrm{cts}}:[0,T]\times\mathscr{X}_{N}\to[0,\infty)$ be the unique solution to (2.9) and let $\alpha_{\mathrm{cts}}(t,x,z)=\hat{\alpha}(x,z,\Delta_{x}v_{\mathrm{cts}}(t,\cdot,z))$. The following proof establishes Theorem 2.23 and shows that $(v^{(K)},\mathfrak{a}^{(K)}/h^{(K)})\to(v_{\mathrm{cts}},\alpha_{\mathrm{cts}})$ uniformly on $[0,T]\times\mathscr{X}_{N}$, as $K\to\infty$.

### v2

@@ 4.2 The mean-field Nash-Lasry-Lions equation
We now discuss (MF-NLL) and first recall that any solution to (MF-NLL), regardless of its regularity in the $\mu$ variable, defines a solution to the MFG system for any initial condition $(t,\mu)\in{\mathcal{T}}\times\Sigma^{d-1}$, and hence gives rise to solutions to the mean-field game, see Proposition 2.19 (ii).
@@
Let Assumptions 2.6 and 2.8 hold. Then, for any Lipschitz continuous function $\varphi:\mathscr{X}\to\mathbb{R}$, there exists a unique map $\hat{\alpha}:\mathscr{X}\to\Sigma^{d-1}$ that satisfies
@@
where $\gamma$ and $L_{\partial\ell}$ are constants from Assumptions 2.8 and 2.20, respectively.
@@
${\mathcal{S}}:=\{\alpha:{\mathcal{X}}\to\Sigma^{d-1}\,:\,\alpha_{y}(x)=0\ \forall y\notin S(x)\},\quad d_{\mathcal{S}}(\alpha,\tilde{\alpha}):=\max_{x\in{\mathcal{X}}}\sum_{y\in S(x)}|\alpha_{y}(x)-\tilde{\alpha}_{y}(x)|,\quad\alpha,\tilde{\alpha}\in{\mathcal{S}}.$
@@
We first note that, if $\mu\in\Sigma^{d-1}$ is a row vector and $A\in\mathbb{R}^{d\times d}$, then $|\mu\cdot A|_{1}\leq\max_{1\leq i\leq d}|A_{i,:}|_{1}$. Define a function $f:\mathbb{R}^{d}\to[0,\infty]$ by $f(a):=\ell(x,\mu,a/h)h$ for $a\in[0,\infty)^{d}$ and $\infty$ otherwise and set $C:=A(h,x)$. Using Assumption 2.8, we again see $f$ satisfies the conditions of Lemma 3.3. Observe that
$\Phi_{x}(\alpha)=\underset{a\in C}{\mathrm{arg\,max}}\,\Bigl\{\sum_{y\in S(x)}\!a_{y}\,(-\varphi(y,\mu\cdot\alpha))-f(a)\Bigr\}.$
Analogously to the proof of Proposition 3.4, Lemma 3.3 implies
$|\Phi_{x}(\alpha)-\Phi_{x}(\tilde{\alpha})|_{1}\leq\frac{h}{2\gamma}\sum_{y\in S(x)}|\varphi(y,\mu\cdot\alpha)-\varphi(y,\mu\cdot\tilde{\alpha})|\leq\frac{hL_{\varphi}}{2\gamma}\,|\mu\cdot(\alpha-\tilde{\alpha})|_{1}\leq\frac{hL_{\varphi}}{2\gamma}\,\max_{x\in{\mathcal{X}}}|\alpha(x)-\tilde{\alpha}(x)|_{1},$
so that $\Phi$ is a contraction since $h<\gamma/L_{\varphi}$, where $L_{\varphi}=m\max_{x\in{\mathcal{X}}}\mathrm{Lip}_{1}(\varphi(x,\cdot))$. By Banach’s fixed-point theorem, there exists a unique fixed point in $\mathcal{S}$}. Since $\mu\in\Sigma^{d-1}$ was arbitrary, this construction yields a unique fixed point $\hat{\alpha}(x,\mu)$. The Lipschitz constant of $\hat{\alpha}$ can be deduced similarly to the finite-player case, see the proof of Proposition 3.4. ∎
@@ 4.3 One-step example
${\mathcal{X}}=\{0,1\},\quad\tilde{A}(h)=[\sigma^{2}h,1-\sigma^{2}h],\quad\ell(x,a)=\frac{1}{2}a^{2},\quad g(x,\mu)=\mu\,\chi_{x=0}+(1-\mu)\,\chi_{x=1},$
for $(x,a,\mu)\in{\mathcal{X}}\times\tilde{A}(h)\times[0,1]$. We assume $\sigma^{2}h\leq 1/2$, so that $\tilde{A}(h)\neq\emptyset$. As in the two-player game in Example 3.6, addition on $\mathcal{X}$} is defined modulo $2$ and we identify controls $\alpha(x)=(\alpha_{0}(x),\alpha_{1}(x))$ with the probability $\alpha_{x+1}(x)\in\tilde{A}(h)$ of changing the current state. Similarly, we only keep track of the fraction $\mu\in[0,1]$ of the population in state $1$. Given a flow $\mu=(\mu_{0},\mu_{h})\in[0,1]^{2}$ of probabilities representing the fractions of players in state $1$, a typical player solves
@@
for every $x\in{\mathcal{X}}$. The cost functional is chosen in a way that leads to synchronization with the population. Given the initial distribution $\mu_{0}$, mean-field game equilibria are characterized by the following system:
$\begin{cases}\alpha(x)=\underset{a\in A(h)}{\mathrm{arg\,min}}\,\{a^{2}/(2h)+(1-a)g(x,\mu_{h})+ag(x+1,\mu_{h})\},\quad x\in{\mathcal{X}},\\ \mu_{h}=\alpha(0)(1-\mu_{0})+(1-\alpha(1))\mu_{0}.\end{cases}$
@@
with $\alpha^{\mu_{h}}(0):=[h(2\mu_{h}-1)]\lor(\sigma^{2}h)\land(1-\sigma^{2}h)$ and $\alpha^{\mu_{h}}(1):=[h(1-2\mu_{h})]\lor(\sigma^{2}h)\land(1-\sigma^{2}h)$. This is a fixed-point equation for the Nash equilibrium distribution $\mu_{h}\in[0,1]$. If $\sigma^{2}=0$ and $h\geq 1$, then there are three equilibria, $\mu_{h}\in\{0,1/2,1\}$, starting from the uniform distribution. If $\sigma^{2}>0$, we have the following result, which follows by inspection of the above fixed-point equation, see Figure 1 for an illustration.
@@
    If $h\in\{h_{*},h^{*}\}$, then there are exactly three NE, including $\mu_{h}=1/2.$
@@
    If $h\in(h_{*},h^{*})$, then there are exactly five NE, including $\mu_{h}=1/2.$
@@
Figure 1: Nash equilibria starting from $\mu_{0}=1/2$ when $0<\sigma^{2}<3-2\sqrt{2}$.
@@ 5 Short-time results
Theorem 2.11 establishes the existence and uniqueness of the solution to ($N$-NLL) for a time step $h<h_{*}(N,T)$ that depends on both the number of players and the time horizon of the game. In fact, as stated in Theorem 2.21, there exist time steps that guarantee uniqueness for any finite number of players as well as for the mean-field limit. We proceed to establish this result.
@@
    (Finite-player case). Let $\varphi:\mathscr{X}_{N}\to\mathbb{R}$ be given. Let $\hat{\alpha}$ be as in Proposition 3.4, and define ψ(x,z) = H(x,z,E(x,z,φ,^α),^α(x,z)),  (x,z)X_N.
@@
    (Mean-field case). Let $\varphi:\mathscr{X}\to\mathbb{R}$ be given. Let $\hat{\alpha}$ be as in Proposition 4.1, and define ψ(x,μ)= H(x,μ,φ(⋅^α(,μ)),^α(x,z)),  (x,μ)X.
@@
[…]
@@
(b) For $\mu,\tilde{\mu}\in\Sigma^{d-1}$ and $a\in A(h,x)$,
[…]
@@
[…]
@@
Let Assumptions 2.6, 2.8 and 2.20 hold. There exist constants $h_{*},T_{*}>0$ depending only on the cost functions $\ell,g$ and $m$, such that, for all $h>0$ and $K\geq 1$, the following holds: If $h<h_{*}$ and $Kh<T_{*}$, then there exists a unique solution $(v^{(\infty)},\alpha^{(\infty)})$ to (MF-NLL) and a unique solution $(v^{(N)},\alpha^{(N)})$ to ($N$-NLL) on $\bar{\mathcal{T}}=\{0,\ldots,Kh\}$ for every $N\geq 1$. Furthermore,
$\sup_{N\in\mathbb{N}\cup\{\infty\}}\max_{(t,x)\in\bar{\mathcal{T}}\times{\mathcal{X}}}\,\big[\mathrm{Lip}_{1}(v^{(N)}(t,x,\cdot))+\mathrm{Lip}_{1}(\alpha^{(N)}(t,x,\cdot))\big]<\infty.$ (5.2)
@@
$L(K)$ $=L(0)+\sum_{k=1}^{K}(L(k)-L(k-1))$
$\leq L(0)+h\sum_{k=1}^{K}\Bigl(mL_{\ell}+L(k-1)\frac{L_{\partial\ell}+L(k-1)}{\gamma-hL(k-1)}\Bigr)\leq L(0)+Kh\,\tilde{M}\leq M,$
@@
Step 2. In light of the previous step, we may choose any $M>L(0)$ and define $h_{*}:=\gamma/M$ and $T_{*}:=(M-L(0))/\tilde{M}$. Note that, in particular, we have $h<\gamma/L(k)$ for all $k=0,\ldots,K-1$. By Propositions 3.4 and 4.1, we can recursively solve the equations ($N$-NLL) and (MF-NLL) on $\{0,\ldots,Kh\}$. Finally, note that the Lipschitz constants of the value functions are bounded by $M$, which implies a bound for the Lipschitz constants of the optimal controls as well. ∎
@@ 6 Mean-field convergence
$\mathbb{E}[\varphi^{(N)}(x,Z^{\prime}(x,z^{(N)},\alpha^{(N)}))]\to\varphi^{(\infty)}\Big(x,\sum_{y\in{\mathcal{X}}}\mu(y)\alpha^{(\infty)}(y,\mu)\Big),\quad\text{as}\quad N\to\infty.$
@@
[…]
@@
$\tilde{\alpha}^{(\infty)}(t,x,\mu)=\underset{a\in A(h,x)}{\mathrm{arg\,min}}\,H(x,\mu,\tilde{v}^{(\infty)}(t+h,\cdot,\mu\cdot\tilde{\alpha}^{(\infty)}(t,\cdot,\mu)),a),\quad(t,x,\mu)\in{\mathcal{T}}\times\mathscr{X}.$
@@
For a given $\mu\in\Sigma^{d-1}$, choose $z^{(N)}\in\mathscr{X}_{N}$ such that $z^{(N)}\to\mu$ as $N\to\infty$. By Lemma 6.1 and continuity of $\ell$,
$H(x,\mu,\tilde{v}^{(\infty)}(t+h,\cdot,\mu\cdot\tilde{\alpha}^{(\infty)}(t,\cdot,\mu)),\tilde{\alpha}^{(\infty)}(t,x,\mu))\leq H(x,\mu,\tilde{v}^{(\infty)}(t+h,\cdot,\mu\cdot\tilde{\alpha}^{(\infty)}(t,\cdot,\mu)),a)$
[…]
@@ 7 Continuous-time limit
This section proves Theorem 2.23. Recall the definitions of $S(x)$ and $A(h,x)$ in (2.11). We again set $m:=\max_{x\in{\mathcal{X}}}|S(x)|$.
@@
Let $(\alpha^{(K)},\alpha^{(\infty)}):\mathscr{X}_{N}\to\Sigma^{d-1}\times\mathbb{R}^{d}$, $K\geq 1$, be functions such that $\mathfrak{a}^{(K)}/h^{(K)}\to\alpha^{(\infty)}$ as $K\to\infty$, where $\mathfrak{a}^{(K)}(x,z):=\alpha^{(K)}(x,z)-e_{x}$ for $(x,z)\in\mathscr{X}_{N}$. Then, for all $(x,z,z^{\prime})\in\mathscr{X}_{N}\times\Sigma^{d-1}_{N}$,
@@
and, in fact, $\lim_{K\to\infty}\mathbb{P}[Z^{\prime}(x,z,\alpha^{(K)})=z]=1$.
@@
Fix $(x,z)\in\mathscr{X}_{N}$. For any state $y\in{\mathcal{X}}$ and any next-step distribution $s\in\Sigma^{d-1}_{Nz_{y}}$ of the untagged players in state $y$,
$\mathbb{P}[Z^{\prime}(x,y,z,\alpha^{(K)})=s]=\frac{(Nz_{y})!}{(Ns_{1})!\cdots(Ns_{d})!}\,[1+\mathfrak{a}_{y}^{(K)}(y,z+e_{xy})]^{Ns_{y}}\prod_{w\neq y}[\mathfrak{a}^{(K)}_{w}(y,z+e_{xy})]^{Ns_{w}}.$
By inspection of this formula, we deduce the following:
@@
    if $s=z_{y}e_{y}$, then P[Z’(x,y,z,α^(K))=z_ye_y] = [1+a^(K)_y(y,z+e_xy)]^Nz_y →1, as K→∞, using $\mathfrak{a}^{(K)}_{y}(y,z+e_{xy})\to 0$;
@@
    if there exists a $w\in{\mathcal{X}}\setminus\{y\}$ such that $s=z_{y}e_{y}+e_{wy}$, then 1h(K) P[Z’(x,y,z,α^(K))=z_ye_y+e_wy] = (Nzy)!(Nzy-1)!   [1+a_y^(K)(y,z+e_xy)]^Nz_y-1a(K)w(y,z+exy)h(K), which converges to $Nz_{y}\,\alpha^{(\infty)}_{w}(y,z+e_{xy})$;
@@
    in all other cases, 1h(K)P[Z’(x,y,z,α^(K))=s] = (Nzy)!(Ns1)!⋯(Nsd)!   [1+a_y^(K)(y,z+e_xy)]^Ns_y 1h(K)∏_i≠y[a^(K)_i(y,z+e_xy)]^Ns_i converges to zero as $\mathfrak{a}_{i}^{(K)}(y,z+e_{xy})/h^{(K)}$ stays bounded as $K\to\infty$ for all $i$ and there are at least two states $i\neq y$ present in the product.
Next, let $z^{\prime}\neq z$ in $\Sigma^{d-1}_{N}$ be given and set $\Sigma^{d-1}_{0}:=\{0_{\mathbb{R}^{d}}\}$. We let $\Sigma(z^{\prime})$ denote the set of all $s=(s(1),\ldots,s(d))$ such that $s(j)\in\Sigma^{d-1}_{Nz_{y}}$ and $z^{\prime}=s(1)+\cdots+s(d)$. Then, the claim follows from the previous observation, together with the explicit formula